# LeviTor: 3D Trajectory Oriented Image-to-Video Synthesis

### v1

@@ Abstract
The intuitive nature of drag-based interaction has led to its growing adoption for controlling object trajectories in image-to-video synthesis. Still, existing methods that perform dragging in the 2D space usually face ambiguity when handling out-of-plane movements. In this work, we augment the interaction with a new dimension, i.e., the depth dimension, such that users are allowed to assign a relative depth for each point on the trajectory. That way, our new interaction paradigm not only inherits the convenience from 2D dragging, but facilitates trajectory control in the 3D space, broadening the scope of creativity. We propose a pioneering method for 3D trajectory control in image-to-video synthesis by abstracting object masks into a few cluster points. These points, accompanied by the depth information and the instance information, are finally fed into a video diffusion model as the control signal. Extensive experiments validate the effectiveness of our approach, dubbed LeviTor, in precisely manipulating the object movements when producing photo-realistic videos from static images. Project page: LeviTor.
@@ 1 Introduction
Our method demonstrates superior performance both quantitatively and qualitatively compared to existing approaches. We achieve accurate 3D trajectory control in image-to-video synthesis task where previous baselines fail, producing realistic and consistent object movements. In summary, our contributions are as follows: We introduce LeviTor, a novel method for controlling 3D object trajectories in video synthesis by combining depth information with K-means clustered points without the need for explicit 3D trajectory tracking. We leverage the high-quality SA-V dataset for training, effectively capturing complex object motions and interactions in diverse scenes. We develop a user-friendly inference pipeline that simplifies the input of 3D trajectories, making it accessible to a broader range of users. To the best of our knowledge, this work is the first to introduce 3D object trajectory control in image-to-video synthesis, paving the way for more advanced and accessible video generation techniques.
@@ 2.1 Video Diffusion Models
Diffusion models [37, 14, 38] have demonstrated unprecedented power in video generation. Video Diffusion Models (VDMs) [15] are broadly categorized into Text-to-Video (T2V) and Image-to-Video (I2V) frameworks, aiming to generate video samples from text prompts or image prompts. T2V generation [5, 7, 30, 36, 4, 41, 46, 16, 52, 21, 9, 10] has been extensively studied in recent years, introducing text descriptions to semantically control the content of video generation. Previous works [5, 12, 42, 44, 59, 10] incorporate temporal layers into large pretrained text-to-image (T2I) diffusion models [34]. Subsequent studies [5, 4, 52, 58, 24] have expanded T2V capabilities by utilizing large text-video pairs, achieving improved results. Building upon T2V, I2V synthesis [55, 4, 52, 58, 8, 49, 27] has also been widely explored. Given a still image, I2V aims to animate it into a video clip that retains all visual content from the image and exhibits naturally suggested dynamics. Many recent works, such as SVD [4], VideoCrafter2 [8], CogVideoX [52] and Open-Sora [58] support both T2V and I2V simultaneously. They extend T2V models to I2V using pretrained T2V models and achieve promising results.
Despite producing high-quality videos, these models rely on text or image prompts, limiting fine-grained control and potentially leading to actions misaligned with user intentions. For precise control, some works [17, 18, 43, 61, 50, 29, 31, 56, 48, 62, 19] employ multimodal video sequences as conditions, such as pose [61, 50, 17, 29, 31, 56], depth [48, 13, 62], or sound [19, 25, 26, 39], treating video generation as a video translation task. Although these models achieve precise control, they require per-frame dense control signals, which makes them cumbersome and not user-friendly in real-world applications. Therefore, simpler yet precise control mechanisms are needed. Trajectory-based control offers an effective method for manipulating video generation, combining simplicity with precision.
@@ 2.2 Trajectory Control in Video Generation
Video synthesis via trajectory control has recently gained popularity due to its ability to achieve precise motion control. Early works [11, 1, 2, 3] employed recurrent neural networks or optical flow to guide motion. Methods like TrailBlazer [28] utilize bounding boxes to direct subject motion in video generation. MotionCtrl [45] encodes trajectory coordinates into dense vector maps, and DragNUWA [53] transforms sparse strokes into dense flow spaces; both use these representations as guidance signals. Tora [57] employs a motion variational autoencoder [22] to embed trajectory vectors into the latent space, preserving motion information across frames.
Although these methods facilitate trajectory control, they often lack semantic understanding of entities, making control over video generation less refined. To address this issue, DragAnything [47] combines entity representation extraction with a 2D Gaussian representation to achieve entity-level controllable video generation. TrackGo [60] uses user-provided free-form masks and arrows to define target regions and movement trajectories, serving as precise blueprints for video generation. However, all these methods consider 2D trajectories in image space, leading to ambiguities in real 3D environments. In this paper, we introduce an innovative control signal representation that combines depth information with K-means clustered points from object masks in video, achieving accurate entity-level and 3D trajectory control.
@@ 3.2 Training Pipeline
Given a VOS format video $V\in\mathbb{R}^{L\times H\times W\times 3}$, it provides the ground truth masks of multiple objects in the video, represented as $\{\{M_{i}^{j}\}_{i=1}^{N}\}_{j=1}^{L}$, where $N$ denotes the number of object masks in each frame. For each mask $M_{i}^{j}$, we conduct K-means algorithm to obtain $k$ center points as control signal. Specifically, we first calculate the area ratio of $M_{i}^{j}$ to the entire image and multiply a hyper-parameter $\alpha$ to determine the approximate number of cluster points:
@@ 4.2 Comparison with Other Approaches
Qualitative comparison results are shown in Fig. 6, where we input the same 2D control trajectory to all models. The top two rows of images show the verification results of occlusion control. In this case, we provide our LeviTor with different depth variations: the depth in the first row changes from far to near, while the depth in the second row only moves closer without being closer to the camera than the buildings on street side. The generated results perfectly meet our requirements, with the tornadoes progressing from far to near and gradually getting larger. Meanwhile, tornado in the first row sweeps across the front of the building, while in the second row it just passes behind the building. In contrast, the other two methods can only control the generation through 2D trajectories. It can be observed that DragAnything misinterprets the movement of the tornado as a forward movement of the camera, resulting in a blurry output. On the other hand, DragNUWA correctly understands that the tornado needs to be moved. However, since it lacks consideration of changes in depth, the size of the tornado hardly changes after the movement, which does not comply with physical laws.
Evaluation results on control for forward and backward object movements in relation to the lens are shown as the left-bottom images in Fig. 6. It is clear that 2D trajectory cannot provide depth information, so DragAnything and DragNUWA can only simulate planets motion that conforms to that trajectory, resulting in blurry and uncertain videos. In contrast, LeviTor can generate accurate and clear movements of two planets based on user-specified inputs meanwhile conforming to physical laws.
@@
Quantitative comparison. We evaluate the quantitative results with two input settings: Single-Point and Multi-Points. The setting of Single-Point is consistent with the evaluation of previous work [47], which means that only one point trajectory is selected for each mask as the condition to generate the video. However, as we mentioned in Sec. 3, one single point trajectory cannot represent forward and backward movements of object in relation to the lens or occlusion. Therefore, we also conduct evaluation with the Multi-Points setting, in which we select at most 8 points in each mask and use their trajectories as condition. Tab. 1 shows the quantitative comparison results of our LeviTor with baselines on DAVIS. Using the same SVD as base model, our method achieves a significant advantage in both FID and FVD metrics, thanks to the consideration of 3D trajectory and training on high-quality VOS dataset SA-V. Besides, increasing the number of control trajectories can effectively benefit DragNUWA and our method. This indicates that considering object size changes over time and occlusion is effective. DragAnything is trained using a single trajectory with object mask semantic information in the first frame, thus increasing the number of trajectories doesn’t match the training and the improvement is limited. Our LeviTor performs worse than DragNUWA on the ObjMC metric, which we attribute to the fact that we do not use tracking methods to obtain complete point trajectories and require the generated video to perfectly match these trajectories. Instead, we directly extracted K-means cluster centers of all masks in each frame as control signal to place greater emphasis on the overall motion and temporal changes of objects.
@@ 4.3 Ablation Studies
In this section, we conduct ablations to study how depth points, instance information and the number of control points for inference affect our synthesis results with the Multi-Points setting.
Depth and instance information. Tab. 2 shows the results of training LeviTor without depth or object instance input, which suggest that both depth and instance information are helpful to our model learning. Compared to depth information, object instance is more important because it represents the objects corresponding to different control points. Without this information, model can easily confuse the control points of different objects, leading to blurred and unrealistic results. Depth information of objects is to some extent implicit in the degree of clustering of points, so its impact is relatively small. We also present a qualitative ablation result in Fig. 7, which suggests that without instance or depth information, the model can easily confuse occlusion relationship between objects, resulting in blurry and unrealistic generation results. Specifically, without object instance information, the model estimates control points that do not belong to itself as its own, leading to phenomena such as the fusion of the sword and person, or deformation of the sword. Without depth information, the model confuses the front and back relationships when inpainting an object based on control points, resulting in cases where the body of sword disappears but the tip of sword passes through the person’s head.
Figure 8: Ablation on number of inference control points.
@@ 5 Conclusion
In this paper, we have presented LeviTor, a novel model for implementing 3D object trajectory control in image-to-video synthesis. Taking depth information combined with K-means clustered points as control signal, our approach captures essential 3D attributes without the need of explicit 3D trajectory estimation. Our user-friendly inference pipeline allows users to input 3D trajectories by simply drawing on 2D images and adjusting point depths, making the synthesis process more accessible. For future work, we aim to extend our model by incorporating more advanced video base models capable of capturing deformable objects and intricate dynamics, in order to better handle non-rigid motions. This enhancement will broaden the applicability of our approach to a wider range.
@@
Appendix A Visualization results
We have placed the visualization results generated by LeviTor in the folder of supplementary material. You can open the HTML file within the folder locally to view the video synthesis results. Please refresh the webpage if the GIFs appear to be out of sync.
@@ Appendix C Comparison with Single-point Control
One of our key motivations is to represent 3D motions by utilizing the clustering and dispersion of multiple points within object masks. Another more intuitive idea is whether we can represent 3D motion using 2D trajectories combined with depth information. That is, representing a 3D trajectory through a single 2D trajectory along with changes of depth values input by users. To validate this idea, we use the center point of each object’s mask as a control point and train the model with the value change of that point as the generation condition. We conduct both qualitative and quantitative analysis. Qualitative results in Fig. S2 show that such single-point control can not represent 3D motions well. The first two examples test the representation of occlusion. It can be observed that a single point with depth changes controlling struggles to accurately express occlusion, resulting in the disappearance of the purple light cluster and the deformation and merging of the cars. The third example tests the control of forward and backward movements. Compared to our LeviTor, single-point control is not very sensitive to size changes caused by forward and backward movement. Quantitative results in Tab. S1 also show the advantage of 3D motion representation with clustering and dispersion of multiple points. The ablation study in Tab. 2 of the main text indicates that the value of depth does not significantly affect the quality of the generated results. And results in this section show that 2D trajectories with depth value changes can not represent 3D motions. These conclusions both suggest that in our method, the clustering and dispersion of multiple control points are the key aspects of 3D motion representation, while depth information is generally used for moving objects in 3D space to obtain rendered object masks.
@@ Appendix D Bad Case Analysis
We, in this section, list some bad generation cases for analysis. Results shown in Fig. S3 indicate that our LeviTor has difficulties in reconstructing small faces and generating scenes with large motions. It may also confuse similar parts of objects. For example, in the first row of Fig. S3, the horse faces become blurry while walking, and the movement of their legs is also quite unnatural. Similarly, in Fig. S1, the movement of the person’s feet while running also appears unnatural. In the second row, the elephant’s front leg suddenly turns into a back one, and then a regenerated front leg appears. We attribute this phenomenon to the fact that the underlying video base model Stable Video Diffusion (SVD) [4] we apply is unable to reconstruct small faces and tends to produce artifacts when generating large-scale movements. We are going to enhance our model by integrating more advanced video-based models in the future, hoping to better capture deformable objects and complex dynamics to handle large-scale and non-rigid motions.

### v2

@@ Abstract
The intuitive nature of drag-based interaction has led to its growing adoption for controlling object trajectories in image-to-video synthesis. Still, existing methods that perform dragging in the 2D space usually face ambiguity when handling out-of-plane movements. In this work, we augment the interaction with a new dimension, i.e., the depth dimension, such that users are allowed to assign a relative depth for each point on the trajectory. That way, our new interaction paradigm not only inherits the convenience from 2D dragging, but facilitates trajectory control in the 3D space, broadening the scope of creativity. We propose a pioneering method for 3D trajectory control in image-to-video synthesis by abstracting object masks into a few cluster points. These points, accompanied by the depth information and the instance information, are finally fed into a video diffusion model as the control signal. Extensive experiments validate the effectiveness of our approach, dubbed LeviTor, in precisely manipulating the object movements when producing photo-realistic videos from static images. Our code is available at: https://github.com/ant-research/LeviTor.
@@ 1 Introduction
Our method demonstrates superior performance both quantitatively and qualitatively compared to existing approaches. We achieve accurate 3D trajectory control in image-to-video synthesis task where previous baselines fail. In summary, our contributions are as follows: We introduce LeviTor, a novel method for controlling 3D object trajectories in video synthesis by combining depth information with K-means clustered points without the need for explicit 3D trajectory tracking. We leverage the high-quality SA-V dataset for training, effectively capturing complex object motions and interactions in diverse scenes. We develop a user-friendly inference pipeline that simplifies the input of 3D trajectories, making it accessible to a broader range of users. To the best of our knowledge, this work is the first to introduce 3D object trajectory control in image-to-video synthesis, paving the way for more advanced and accessible video generation techniques.
@@ 2.1 Video Diffusion Models
Diffusion models [43, 17, 44] have demonstrated unprecedented power in video generation. Video Diffusion Models (VDMs) [18] are broadly categorized into Text-to-Video (T2V) and Image-to-Video (I2V) frameworks, aiming to generate video samples from text prompts or image prompts. T2V generation [6, 8, 33, 42, 5, 47, 52, 19, 58, 24, 12, 13] has been extensively studied in recent years, introducing text descriptions to control the content of video generation semantically. Previous works [6, 15, 48, 50, 67, 13] incorporate temporal layers into large pretrained text-to-image (T2I) diffusion models [40]. Subsequent studies [6, 5, 58, 66, 27] have expanded T2V capabilities by utilizing large text-video pairs, achieving improved results. Building upon T2V, I2V synthesis [63, 5, 58, 66, 9, 55, 30] has also been widely explored. Given a still image, I2V aims to animate it into a video clip that retains all visual content from the image and exhibits naturally suggested dynamics. Many recent works, such as SVD [5], VideoCrafter2 [9] and CogVideoX [58] support both T2V and I2V simultaneously.
Despite producing high-quality videos, these models rely on text or image prompts, limiting fine-grained control and potentially leading to actions misaligned with user intentions. For precise control, some works [20, 21, 49, 69, 56, 32, 37, 64, 54, 70, 22, 61] employ multimodal video sequences as conditions, such as pose [69, 56, 20, 32, 37, 64], depth [54, 16, 70], or sound [22, 28, 29, 45], treating video generation as a video translation task. Although these models achieve precise control, they require per-frame dense control signals, which makes them cumbersome and not user-friendly in real-world applications. Therefore, simpler yet precise control mechanisms are needed. Trajectory-based control offers an effective method for manipulating video generation, combining simplicity with precision.
@@ 2.2 Trajectory Control in Video Generation
Controllable editing has gained advancements in the field of image editing due to its precise control information [10, 34, 59, 2]. For video synthesis, trajectory-controlled generation has recently gained popularity due to its ability to achieve precise motion control. Early works [14, 1, 3, 4] employed recurrent neural networks or optical flow to guide motion. Methods like TrailBlazer [31] utilize bounding boxes to direct subject motion in video generation. MotionCtrl [51] encodes trajectory coordinates into dense vector maps, and DragNUWA [60] transforms sparse strokes into dense flow spaces; both use these representations as guidance signals. Tora [65] employs a motion variational autoencoder [25] to embed trajectory vectors into the latent space, preserving motion information across frames.
Although these methods facilitate trajectory control, they often lack semantic understanding of entities, making control over video generation less refined. To address this issue, DragAnything [53] combines entity representation extraction with a 2D Gaussian representation to achieve entity-level controllable video generation. TrackGo [68] uses user-provided free-form masks and arrows to define target regions and movement trajectories, serving as precise blueprints for video generation. However, all these methods consider 2D trajectories in image space, leading to ambiguities in real 3D environments. The recent 3D-TrajMaster [11] manipulates multi-entity 3D motions with user-desired 6DoF pose sequences of entities for video generation. In this paper, we introduce an innovative control signal representation that combines depth information with K-means clustered points from object masks in video, achieving accurate entity-level and 3D trajectory control.
@@ 3.2 Training Pipeline
Given a VOS format video $V\in\mathbb{R}^{L\times H\times W\times 3}$, it provides the ground truth masks of multiple objects in the video, represented as $\{\{M_{i}^{j}\}_{i=1}^{X}\}_{j=1}^{L}$, where $X$ denotes the number of object masks in each frame. For each mask $M_{i}^{j}$, we conduct K-means algorithm to obtain $k$ center points as control signal. Specifically, we first calculate the area ratio of $M_{i}^{j}$ to the entire image and multiply a hyper-parameter $\alpha$ to determine the approximate number of cluster points:
@@ 4.2 Comparison with Other Approaches
Qualitative comparison results are shown in Fig. 6, where we input the same 2D control trajectory to all models. The top two rows of images show the verification results of occlusion control. In this case, we provide our LeviTor with different depth variations: the depth in the first row changes from far to near, while the depth in the second row only moves closer without being closer to the camera than the buildings on street side. The generated results perfectly meet our requirements, with the tornadoes progressing from far to near and gradually getting larger. Meanwhile, tornado in the first row sweeps across the front of the building, while in the second row it just passes behind the building. In contrast, the other two methods can only control the generation through 2D trajectories. It can be observed that DragAnything misinterprets the movement of the tornado as a forward movement of the camera, resulting in a blurry output. On the other hand, DragNUWA correctly understands that the tornado needs to be moved. However, since it lacks consideration of changes in depth, the size of the tornado hardly changes after the movement, which does not comply with perspective projection rules.
Evaluation results on control for forward and backward object movements in relation to the lens are shown as the left-bottom images in Fig. 6. It is clear that 2D trajectory cannot provide depth information, so DragAnything and DragNUWA can only simulate planets motion that conforms to that trajectory, resulting in blurry videos. In contrast, LeviTor can generate accurate and clear movements of two planets based on user-specified inputs meanwhile conforming to perspective projection rules.
@@
Quantitative comparison. We evaluate the quantitative results with two input settings: Single-Point and Multi-Points. The setting of Single-Point is consistent with previous work [53], which means that only one point trajectory is selected for each mask as video generation condition. For Multi-Points setting, we select at most 8 points in each mask and use their trajectories as condition. Tab. 1 shows the quantitative comparison results of LeviTor with baselines on DAVIS. Using the same SVD as base model, our method achieves a significant advantage in both FID and FVD metrics, thanks to the consideration of 3D trajectory and training on high-quality SA-V dataset. Besides, increasing the number of control trajectories can effectively benefit DragNUWA and LeviTor. This indicates that considering object size changes over time and occlusion is effective. DragAnything is trained using a single trajectory with object mask semantic information in first frame, thus increasing the number of trajectories doesn’t match the training and improvement is limited. LeviTor performs worse than DragNUWA on the ObjMC metric, which we attribute to the fact that we do not use tracking methods to obtain complete point trajectories and require the generated video to perfectly match these trajectories.
@@ 4.3 Ablation Studies
We conduct ablations to study how depth points, instance information and the number of control points for inference affect our synthesis results with the Multi-Points setting.
Depth and instance information. Tab. 2 shows the results of training LeviTor without depth or object instance input, which suggest that both depth and instance information are helpful to our model learning. Compared to depth information, object instance is more important because it represents the objects corresponding to different control points. Without this information, model can easily confuse the control points of different objects, leading to blurred and unrealistic results. Depth information of objects is to some extent implicit in the degree of clustering of points, so its impact is relatively small. We also present a qualitative ablation result in Fig. 7, which suggests that without instance or depth information, the model can easily confuse occlusion relationship between objects, resulting in blurry and unrealistic generation results.
Figure 8: Ablation on number of inference control points. ’Scale’ means the value multiplied by the default number of control points.
@@ 5 Conclusion
In this paper, we have presented LeviTor, a novel model for implementing 3D object trajectory control in image-to-video synthesis. Taking depth information combined with K-means clustered points as control signal, our approach captures essential 3D attributes without the need for explicit 3D trajectory estimation. Our user-friendly inference pipeline allows users to input 3D trajectories by simply drawing on 2D images and adjusting point depths, making the synthesis process more accessible. Our model also has certain limitations. First, LeviTor is constrained by the segmentation results of SAM and trajectories provided by the user, and it does not understand physical laws to generate movements of objects without provided trajectory controls. Additionally, since LeviTor was not trained using tracking data, it cannot control the internal pose changes of objects. Finally, the current generation results are limited by the base model SVD. For future work, we aim to extend LeviTor by incorporating more advanced video base models capable of capturing deformable objects and intricate dynamics to better handle non-rigid motions.
Acknowledgements: This work is supported by the National Key R&\&&D Program of China (No. 2022ZD0160900), the Research Grant Council of the Hong Kong Special Administrative Region under grant number 16203122, Ant Group Research Intern Program, Jiangsu Frontier Technology Research and Development Program (No. BF2024076), and the Collaborative Innovation Center of Novel Software Technology and Industrialization.
[…]
Supplementary Material
@@
Appendix A Comparison with more methods
This section compares our LeviTor with more recent methods SG-I2V [35] and MOFA-Video [36]. The qualitative comparison in  Fig. S1 shows that these methods fail to follow complex trajectories or produce proper depth variation.
Figure S1: Qualitative comparison with SG-I2V and MOFA-Video.
@@ Appendix C Comparison with Single-point Control
One of our key motivations is to represent 3D motions by utilizing the clustering and dispersion of multiple points within object masks. Another more intuitive idea is whether we can represent 3D motion using 2D trajectories combined with depth information. That is, representing a 3D trajectory through a single 2D trajectory along with changes of depth values input by users. To validate this idea, we use the center point of each object’s mask as a control point and train the model with the value change of that point as the generation condition. We conduct both qualitative and quantitative analysis. Qualitative results in Fig. S3 show that such single-point control can not represent 3D motions well. The first two examples test the representation of occlusion. It can be observed that a single point with depth changes controlling struggles to accurately express occlusion, resulting in the disappearance of the purple light cluster and the deformation and merging of the cars. The third example tests the control of forward and backward movements. Compared to our LeviTor, single-point control is not very sensitive to size changes caused by forward and backward movement. Quantitative results in Tab. S1 also show the advantage of 3D motion representation with clustering and dispersion of multiple points. Ablation study in Tab. 2 of the main text indicates that the value of depth does not significantly affect the quality of the generated results. And results in this section show that 2D trajectories with depth value changes can not represent 3D motions. These conclusions both suggest that in our method, the clustering and dispersion of multiple control points are the key aspects of 3D motion representation, while depth information is generally used for moving objects in 3D space to obtain rendered object masks.
@@ Appendix D Bad Case Analysis
In this section, we list some bad generation cases for analysis. Results in Fig. S4 indicate that our LeviTor has difficulties in reconstructing small faces and generating scenes with large motions. It may also confuse similar parts of objects. For example, in the first row of Fig. S4, the horse’s faces become blurry while walking, and the movement of their legs is also quite unnatural. In Fig. S2, the movement of the person’s feet while running also appears unnatural. In the second row, the elephant’s front leg suddenly turns into a back one, and then a regenerated front leg appears. We attribute this phenomenon to the fact that the underlying video base model Stable Video Diffusion (SVD) [5] we apply can not reconstruct small faces and tends to produce artifacts when generating large-scale movements. We are going to enhance our model by integrating more advanced video-based models in the future, hoping to better capture deformable objects and complex dynamics to handle large-scale and non-rigid motions.